# ConsistI2V: Enhancing Visual Consistency for Image-to-Video Generation

### v1

@@ Appendix C I2V-Bench Evaluation Metrics
We measure the temporal consistency of the background scenes by computing the CLIP (Radford et al., 2021) feature similarity across frames.
@@
We measure the subject consistency by calculating the DINO (Caron et al., 2021) feature similarity across frames.
@@
We calculate the average absolute difference between each frame.
@@
We adapt AMT (Li et al., 2023) to evaluate the level of smoothness in the generated motions.
@@
We adapt RAFT (Teed & Deng, 2020) to estimate the degree of dynamics in the video.
@@
We medially select 100 videos separately from the Pet, Vehicle, Animal, and Food categories in the I2V-Bench validation dataset. We further annotate the related objects based on the reference videos and use GRiT (Wu et al., 2022) to assess the success rate of generating the stated objects matching the reference videos.
@@
We select 100 videos from the Scenery-Nature and Scenery-City categories in the I2V-Bench validation dataset. We further annotate the related scenes based on the reference videos and apply Tag2Text (Huang et al., 2023a) to caption the generated videos and detect whether the scenes, in the reference videos, exist in them.
@@
We adopt the overall video-text consistency computed by ViCLIP (Wang et al., 2023e) to reflect the semantics consistency of the manually annotated captions of I2V-Bench and the generated videos.
@@ Appendix D Human Evaluation Details
Appendix E Limitations
Appendix F I2V-Bench Statistics
@@ Appendix G Additional Quantitative Results
Appendix H I2V-Bench Results
Table 6: Appendix: Results of I2V-Bench for Background Consistency. S-N, S-C, A-H and A-S respectively represent Scenery-Nature, Scenery-City, Animation-Hard and Animation-Static.
@@ Appendix H I2V-Bench Results
Table 7: Appendix: Results of I2V-Bench for Subject Consistency. S-N, S-C, A-H and A-S respectively represent Scenery-Nature, Scenery-City, Animation-Hard and Animation-Static.
@@
Table 8: Appendix: Results of I2V-Bench for Temporal Flickering. S-N, S-C, A-H and A-S respectively represent Scenery-Nature, Scenery-City, Animation-Hard and Animation-Static.
@@
Table 9: Appendix: Results of I2V-Bench for Motion Smoothness. S-N, S-C, A-H and A-S respectively represent Scenery-Nature, Scenery-City, Animation-Hard and Animation-Static.
@@
Table 10: Appendix: Results of I2V-Bench for object consistency.
@@
Table 11: Appendix: Results of I2V-Bench for scene consistency. S-N and S-C respectively represent Scenery-Nature and Scenery-City
@@
Table 12: Appendix: Results of I2V-Bench for Dynamic Degree. S-N, S-C, A-H and A-S respectively represent Scenery-Nature, Scenery-City, Animation-Hard and Animation-Static.
@@
Table 13: Appendix: Results of I2V-Bench for Text-Video Overall Consistency. S-N, S-C, A-H and A-S respectively represent Scenery-Nature, Scenery-City, Animation-Hard and Animation-Static.

### v2

@@ Appendix C I2V-Bench Evaluation Metrics
We measure the temporal consistency of the background scenes by computing the CLIP (Radford et al., 2021) feature similarity across frames. CLIP features represent the high-level semantic content of an image and are not sensitive to subject variations, making it a suitable feature to measure background consistency.
@@
We measure the subject consistency by calculating the DINO (Caron et al., 2021) feature similarity across frames. Since DINO features capture fine-grained semantic information in the foreground subjects, they are suitable for evaluating subject consistency.
@@
We calculate the average absolute difference between each frame. Temporal flickerings in generated videos will cause large local pixel value variations, thereby causing large frame differences.
@@
We adapt AMT (Li et al., 2023) to evaluate the level of smoothness in the generated motion. Specifically, given a generated video, we drop the odd-number frames to obtain a video with a lower frame rate. We then use AMT to interpolate the video back to the original frame rate. Motion smoothness is defined as the mean absolute error (MAE) between the reconstructed frames and the dropped frames. A large MAE indicates that the dropped frames contain abrupt motions that are not likely to be interpolated by AMT, thus indicating a lack of smooth motion progression in the generated video.
@@
We adopt RAFT (Teed & Deng, 2020) to estimate the degree of dynamics in the video. RAFT is a method designed for optical flow estimation. We use the magnitude of the flow vector to indicate the degree of dynamics presented in the video.
@@
We medially select 100 videos from the Pet, Vehicle, Animal, and Food categories in the I2V-Bench validation dataset. We then annotate each video based on the objects presented in the video (e.g. “cat”, “car”, etc.) and use GRiT (Wu et al., 2022) to determine if the generated video reflects the annotated objects. Object consistency is computed as the success rate of synthesizing the annotated objects in the generated videos.
@@
We select 100 videos from the Scenery-Nature and Scenery-City categories in the I2V-Bench validation dataset. We then annotate the scenes in the videos as different keywords (e.g. “ocean”, “sky”). Scene consistency is computed by applying Tag2Text (Huang et al., 2023a) to caption the generated videos and detect whether the captions in the generated videos contain the annotated keywords.
@@
We adopt the overall video-text consistency by computing the cosine similarity between video and text embeddings using ViCLIP (Wang et al., 2023e) to reflect the semantics consistency of the manually annotated captions of I2V-Bench and the generated videos. ViCLIP is a CLIP-based model trained on video-text data that maps videos and text in a joint embedding space. Therefore, this similarity represents an overall degree of alignment between the generated video and the text prompt.
@@ Appendix D Human Evaluation Details
Appendix E I2V-Bench Statistics
@@ Appendix F Additional Quantitative Results
Appendix G I2V-Bench Results
We present the detailed breakdown of I2V-Bench results in Table 6789101112 and 13.
Table 6: Results of I2V-Bench for Background Consistency. S-N, S-C, A-H and A-S respectively represent Scenery-Nature, Scenery-City, Animation-Hard and Animation-Static.
@@ Appendix G I2V-Bench Results
Table 7: Results of I2V-Bench for Subject Consistency. S-N, S-C, A-H and A-S respectively represent Scenery-Nature, Scenery-City, Animation-Hard and Animation-Static.
@@
Table 8: Results of I2V-Bench for Temporal Flickering. S-N, S-C, A-H and A-S respectively represent Scenery-Nature, Scenery-City, Animation-Hard and Animation-Static.
@@
Table 9: Results of I2V-Bench for Motion Smoothness. S-N, S-C, A-H and A-S respectively represent Scenery-Nature, Scenery-City, Animation-Hard and Animation-Static.
@@
Table 10: Results of I2V-Bench for object consistency.
@@
Table 11: Results of I2V-Bench for scene consistency. S-N and S-C respectively represent Scenery-Nature and Scenery-City
@@
Table 12: Results of I2V-Bench for Dynamic Degree. S-N, S-C, A-H and A-S respectively represent Scenery-Nature, Scenery-City, Animation-Hard and Animation-Static.
@@
Table 13: Results of I2V-Bench for Text-Video Overall Consistency. S-N, S-C, A-H and A-S respectively represent Scenery-Nature, Scenery-City, Animation-Hard and Animation-Static.